# PyTomography: A Python Library for Medical Image Reconstruction

### v5

@@ 2.1.3 Reconstruction Algorithms
$f^{n+1}=f^{n}+C^{n}(f^{n},H^{n})\left[\nabla_{f}L(g^{n}|f^{n},H^{n})-\beta^{n}\nabla_{f}V(f^{n})\right]$ (5)
where $n$ is the subiteration index (used as a superscript and not an exponent for all variables), $f^{n}$ is the object estimate of subiteration $n$, $g^{n}$ specifies a subset of data used in subiteration $n$, $H^{n}$ is a system matrix projecting only to the current subset, $C^{n}(f^{n})$ is a preconditioner factor that depends on the object estimate, $V(f^{n})$ is a prior function used to regularize the algorithm, and $\beta^{n}=c^{n}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c^{n}$). These algorithms partition the data $g$ into $M$ approximately equal subsets; a single iteration of an algorithm is defined as the number of subiterations where all subsets have been used once. Various PGA algorithms available in PyTomography are shown in Table 2; they are characterized by their preconditioner factor.
@@
Table 2: Sample PyTomography classes and their corresponding identifying features. $s$ corresponds to additive term in likelihoods. The sample algorithms shown are the ordered subset expectation maximum (OSEM) [24], ordered subset maximum a posteriori one step late (OSMAPOSL) [25], block sequential regularized expectation maximum (BSREM) [26]. $n$ is used as a superscript (not an exponent) for all variables. For BSREM, $\alpha^{n}$ is a relaxation parameter. The sample nearest neighbour prior functions are characterized by their image weighting function $\phi(f_{r},f_{s})$.
@@
Algorithm classes Preconditioner factor $C^{n}(f^{n},H^{n})$
OSEM $f^{n}/\left(\left(H^{n}\right)^{T}1\right)$ and $V(f^{n})=0$
OSMAPOSL $f^{n}/\left(\left(H^{n}\right)^{T}1+\nabla_{f}V(f^{n})\right)$
BSREM $f^{n}\alpha^{n}/\left(c^{n}H^{T}1\right)$
@@ 2.2 Software architecture
Sample code demonstrating the modularity of the library is shown in Figure 1 for a simplified SPECT use case (e.g. no attenuation, PSF, or scatter correction). Lines 1-9 demonstrate reading the data/metadata, lines 10-13 demonstrate constructing the projector using the loaded data/metadata, lines 14-17 demonstrate creating the likelihood using the system matrix and loaded data, and lines 18-22 demonstrate creating a reconstruction algorithm and obtaining a reconstructed image. A new contribution to the library might involve adding a new system model and replacing lines 1-13 while reusing lines 14-22. Another possible contribution might include a new reconstruction algorithm; in this case lines 18-22 would need to be replaced while the SPECT projector code can be reused.
@@ 3.2 PET
The simulated scanner consisted of $8$ rings of 338 mm radius with 32.25 mm ring spacing, and a TOF coincidence resolution of 550 ps; this fictional scanner is representative of a Siemens Biograph mMR with TOF capabilities. The digital phantom used for simulation in GATE consisted of an ultra-high resolution 18F-FDG PET image, a T1-weighted MRI brain image, and a corresponding attenuation map [28]. The total uptake of 15.66 MBq was representative of a typical 18F-FDG distribution 1 hour post injection. The normalization $\eta$ was obtained via simulation of a thin cylindrical shell of radius 318318318~{}318mm that extended along the axial FOV of the scanner; normalization computation is available in PyTomography for basic polygonal PET geometries. In all cases, simulated data were reconstructed using a $204\times 204\times 154$ matrix size with voxel spacing of 1.251.251.25~{}1.25mm; 5 TOF bins were used with a total range of 30 cm. Before reconstruction, the “randoms” contribution to each listmode event was estimated from delayed coincidences; this process consisted of (i) binning the delayed coincidence events into a sinogram (ii) smoothing the sinogram to reduce noise, and (iii) converting back into LM form for use in LM reconstruction. The “scatters” contribution for each listmode event was estimated using the TOF single scatter simulation (SSS) capabilities of the library, based on Watson [29]. The PET image used as a proxy for SSS was a reconstruction of the data via LM-TOF OSEM (40it, 1ss) without scatter estimation. Simulation and reconstruction were then performed in the following three ways:
@@
    A 9 minute acquisition (high count) was simulated in GATE that included all relevant physical phenomenon in a typical PET scan. The data was then reconstructed using LM-TOF OSEM (4it, 20ss) with attenuation, random, and scatter modeling. PSF modeling was implemented via 3 dimensional blurring in object space using a Gaussian kernel with a full width half max of 3 mm.
@@
    The dataset above was subsampled into a 3 minute acquisition (low count), and an identical reconstruction was performed except with 80it and 1ss.
@@
Figure 3: Top: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [28]. From top to bottom: axial, coronal, and sagittal slices. From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (4it/20ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction. Bottom: Variation in absolute bias and standard deviation throughout iterative reconstruction. Subplots correspond to the grey matter region (left) and white matter region (right). For the low count rate scenario, points are plotted every 10 iterations, and for the low count OSEM example, the first point corresponds to the 10th iteration. The bias was initially positive and became negative for all three cases in white matter.
Table 3: Computational times for all use cases considered in this paper. Times required to open raw data are not considered. Times to construct SPECT system matrices and compute PET randoms in PyTomography are not considered since these are negligible. The notation “s / it” corresponds to the seconds per iteration for a given reconstruction algorithm, ‘SSS’ refers to the single scatter simulation in PET, and ‘SM’ refers to the creation of the system matrix. For the DIPRecon PET case, ‘LC’ refers to the prior low count reconstruction, since this is required to obtain an initial image for DIPRecon.
SPECT DICOM Time Required
@@
SPECT SIMIND Time Required
@@
PET GATE Time Required
High Count OSEM (4it/20ss) $(16.9~{}\text{s/it})&\times 4\text{it}+37.4~{}\text{s (SM)}\\ +&343.8~{}\text{s (SSS)}=\textbf{449.0~{}\text{s}}\end{aligned}$
Low Count OSEM (80it/1ss) $(3.15~{}\text{s/it})&\times 80\text{it}+30.8~{}\text{s (SM)}\\ +&375.7~{}\text{s (SSS)}=\textbf{658.3~{}\text{s}}\end{aligned}$
Low Count DIPRecon (100it) $(21.5~{}\text{s/it})&\times 100\text{it}+658.3~{}\text{s (LC)}\\ =&\textbf{2808.5~{}\text{s}}\end{aligned}$
SPECT Validation Time Required
@@
PET Validation Time Required
@@ 4 Impact
A variety of algorithms were examined for 177Lu SPECT reconstruction; various quantitative metrics were estimated to highlight the differences between the available algorithms in the library. For both the simulated and clinical data, the algorithms yielding smallest to largest kidney concentration were BSREM-RDP, OSEM, KEM, and BSREM-RDPAP. The simulated data prediction from BSREM-RDPAP (802.31802.31802.31~{}802.31kBq/mL) was closest to the ground truth (843.36843.36843.36~{}843.36kBq/mL), suggesting that algorithms incorporating anatomical information may yield the most accurate kidney activity concentration. For both the simulated and clinical data, the order of algorithms yielding the lowest to highest noise were BSREM-RDP, BSREM-RDPAP, OSEM, and KEM; this is unsurprising as the RDP prior is designed to reduce noise. While BSREM-RDPAP has a lower noise than OSEM in each case, it also had higher uptake in the kidney: BSREM-RDPAP thus has simultaneous improvements in both bias and noise reduction. For the bone lesion, algorithms incorporating anatomical information (BSREM-RDPAP and KEM) were able to contain activity mainly inside the bone; this is demonstrated in the axial slices shown in Figure 2. These algorithms also yielded higher bone lesion activity concentration predictions than OSEM for both the simulated/clinical data, with the simulated data concentrations (347.77347.77347.77~{}347.77kBq/mL for BSREM-RDPAP and 220.56220.56220.56~{}220.56kBq/mL for KEM) closer to the ground truth (819.79819.79819.79~{}819.79kBq/mL) than OSEM (205.21205.21205.21~{}205.21kBq/mL). For the liver lesion in the simulated data, however, use of BSREM-RDPAP actually yielded a smaller activity concentration prediction (261.32261.32261.32~{}261.32kBq/mL) compared to OSEM (267.47267.47267.47~{}267.47kBq/mL). This occured because there was no anatomical difference in the surrounding tissue in the liver metastasis, while there was for the bone metastasis. These results, which are consistent with previous studies [6], suggest that anatomical priors may yield more accurate activity concentrations for bone lesion dosimetry. While this example focused primarily on organ and lesion dosimetry, PyTomography users will be to assess the merits of various reconstruction algorithms on any task specific metric.
@@ Appendix A Validation of PyTomography
    Monte Carlo SPECT Validation With STIR: Radioactivity concentrations in a digital phantom were selected to correspond to a 24-hour post injection prostate cancer patient (1700 MBq total activity) imaged via the prostate-specific membrane antigen (PSMA)-targeting radiopharmaceutical, 177Lu-PSMA-617. SPECT acquisition was simulated in SIMIND with dimensions $128\times 384$, a pixel size of $0.48~{}\text{cm}\times 0.48~{}\text{cm}$, 120 projections, an acquisition time of 15 s/projection, and use of Siemens medium energy collimators. Data were reconstructed with OSEM (2 iterations, 8 subsets) using (i) STIR and (ii) PyTomography. Coronal slices and 1D profiles of reconstructions are shown in Figure 4 and are nearly identical. The image noise, quantified by the standard deviation of predicted counts in the liver, was 0.789 counts (STIR) and 0.793 counts (PyTomography). The mean predicted counts in the kidneys were 1.764 (STIR) and 1.761 (PyTomography). The reconstruction time in PyTomography (5.4 s) was significantly faster than the time required in STIR (51542.2 s), owing to the speed-up factor of GPU implementation.
@@
Figure 6: Sample coronal slices (top) and 1D profiles (bottom) of the high resolution brain phantom reconstructed using CASToR and PyTomography with LM TOF OSEM (2it/21ss). The two images were normalized such that the sum of their voxels was identical. 1D profiles correspond to the horizontal lines shown on the coronal slices

### v6

@@ 2.1.3 Reconstruction Algorithms
$f^{(n+1)}=f^{(n)}+C^{(n)}(f^{(n)},H^{(n)})\left[\nabla_{f}L(g^{(n)}|f^{(n)},H^{(n)})-\beta^{(n)}\nabla_{f}V(f^{(n)})\right]$ (5)
where $n$ is the subiteration index (used as a superscript and not an exponent for all variables), $f^{(n)}$ is the object estimate of subiteration $n$, $g^{(n)}$ specifies a subset of data used in subiteration $n$, $H^{(n)}$ is a system matrix projecting only to the current subset, $C^{(n)}(f^{(n)})$ is a preconditioner factor that depends on the object estimate, $V(f^{(n)})$ is a prior function used to regularize the algorithm, and $\beta^{(n)}=c^{(n)}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c^{(n)}$). These algorithms partition the data $g$ into $M$ approximately equal subsets; a single iteration of an algorithm is defined as the number of subiterations where all subsets have been used once. Various PGA algorithms available in PyTomography are shown in Table 2; they are characterized by their preconditioner factor.
@@
Table 2: Sample PyTomography classes and their corresponding identifying features. $s$ corresponds to additive term in likelihoods. The sample algorithms shown are the ordered subset expectation maximum (OSEM) [24], ordered subset maximum a posteriori one step late (OSMAPOSL) [25], block sequential regularized expectation maximum (BSREM) [26]. $n$ is used as a superscript (not an exponent) for all variables. For BSREM, $\alpha^{(n)}$ is a relaxation parameter. The sample nearest neighbour prior functions are characterized by their image weighting function $\phi(f_{r},f_{s})$.
@@
Algorithm classes Preconditioner factor $C^{(n)}(f^{(n)},H^{(n)})$
OSEM $f^{(n)}/\left(\left(H^{(n)}\right)^{T}1\right)$ and $V(f^{(n)})=0$
OSMAPOSL $f^{(n)}/\left(\left(H^{(n)}\right)^{T}1+\nabla_{f}V(f^{(n)})\right)$
BSREM $f^{(n)}\alpha^{(n)}/\left(c^{(n)}H^{T}1\right)$
@@ 2.2 Software architecture
Figure 1 shows a graphical depiction of the structure of the library. Sample code demonstrating the modularity of the library is shown in Figure 2 for a simplified SPECT use case (e.g. no attenuation, PSF, or scatter correction). Lines 1-9 demonstrate reading the data/metadata, lines 10-13 demonstrate constructing the projector using the loaded data/metadata, lines 14-17 demonstrate creating the likelihood using the system matrix and loaded data, and lines 18-22 demonstrate creating a reconstruction algorithm and obtaining a reconstructed image. A new contribution to the library might involve adding a new system model and replacing lines 1-13 while reusing lines 14-22. Another possible contribution might include a new reconstruction algorithm; in this case lines 18-22 would need to be replaced while the SPECT projector code can be reused.
Figure 1: Software architecture diagram for PyTomography. Metadata and projectors consist of an independent component of the library that can be used for analytical system modeling, but can also be used to construct likelihoods that can be passed to image reconstruction algorithms. Implementations of various components displayed here can be found in Table 2.
@@ 3.2 PET
The simulated scanner consisted of $8$ rings of 338 mm radius with 32.25 mm ring spacing, and a TOF coincidence resolution of 550 ps; this fictional scanner is representative of a Siemens Biograph mMR with TOF capabilities. The digital phantom used for simulation in GATE consisted of an ultra-high resolution 18F-FDG PET image, a T1-weighted MRI brain image, and a corresponding attenuation map [28]. The total uptake of 15.66 MBq was representative of a typical 18F-FDG distribution 1 hour post injection. The normalization $\eta$ was obtained via simulation of a thin cylindrical shell of radius 318318318~{}318mm that extended along the axial FOV of the scanner; normalization computation is available in PyTomography for basic polygonal PET geometries. In all cases, simulated data were reconstructed using a $204\times 204\times 154$ matrix size with voxel spacing of 1.251.251.25~{}1.25mm; 21 TOF bins were used with a total range of 30 cm. Before reconstruction, the “randoms” contribution to each listmode event was estimated from delayed coincidences; this process consisted of (i) binning the delayed coincidence events into a sinogram (ii) smoothing the sinogram to reduce noise, and (iii) converting back into LM form for use in LM reconstruction. The “scatters” contribution for each listmode event was estimated using the TOF single scatter simulation (SSS) capabilities of the library, based on Watson [29]. The PET image used as a proxy for SSS was a reconstruction of the data via LM-TOF OSEM (40it, 1ss) without scatter estimation. Simulation and reconstruction were then performed in the following three ways:
@@
    A 9 minute acquisition (high count) was simulated in GATE that included all relevant physical phenomenon in a typical PET scan. The data was then reconstructed using LM-TOF OSEM for up to 27 iterations with 3 subsets with attenuation, random, and scatter modeling. PSF modeling was implemented via 3 dimensional blurring in object space using a Gaussian kernel with a full width half max of 3 mm.
@@
    The dataset above was subsampled into a 3 minute acquisition (low count), and an identical reconstruction was performed for up to 81 iterations with 1 subset.
@@
Figure 4: Top: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [28]. From top to bottom: axial, coronal, and sagittal slices. From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (18it/3ss shown), low count OSEM reconstruction (54it/1ss shown), and low count DIP reconstruction. Bottom: Variation in absolute bias and standard deviation throughout iterative reconstruction; points are plotted at the iteration corresponding to the displayed images. Subplots correspond to the grey matter region (left) and white matter region (right). The bias was initially positive and became negative for all three cases in white matter.
Table 3: Computational times for all use cases considered in this paper. Times required to open raw data are not considered. Times to construct SPECT system matrices and compute PET randoms in PyTomography are not considered since these are negligible. The notation “s / it” corresponds to the seconds per iteration for a given reconstruction algorithm, ‘SSS’ refers to the single scatter simulation in PET, and ‘SM’ refers to the creation of the system matrix. For the DIPRecon PET case, ‘IF’ refers to the initial model fit required for the DIPRecon. All timing performed on the same GPU-optimized system except for the CASToR PET validation example, which was ran on a CPU-optimized system.
SPECT DICOM Computation Time
@@
SPECT SIMIND Computation Time
@@
PET GATE Computation Time
High Count OSEM (27it/3ss) $(21.1~{}\text{s/it})&\times 27\text{it}+30.9~{}\text{s (SM)}\\ +&925.9~{}\text{s (SSS)}=\textbf{1525.5~{}\text{s}}\end{aligned}$
Low Count OSEM (80it/1ss) $(6.69~{}\text{s/it})&\times 80\text{it}+30.8~{}\text{s (SM)}\\ +&1023.7~{}\text{s (SSS)}=\textbf{1589.8~{}\text{s}}\end{aligned}$
Low Count DIPRecon (100it) $(25.5~{}\text{s/it})&\times 100\text{it}+663.8~{}\text{s (IF)}\\ =&\textbf{3209.6~{}\text{s}}\end{aligned}$
SPECT Validation Computation Time
@@
PET Validation Computation Time
@@ 4 Impact
A variety of algorithms were examined for 177Lu SPECT reconstruction; various quantitative metrics were estimated to highlight the differences between the available algorithms in the library. For both the simulated and clinical data, the algorithms yielding smallest to largest kidney concentration were BSREM-RDP, OSEM, KEM, and BSREM-RDPAP. The simulated data prediction from BSREM-RDPAP (807.88807.88807.88~{}807.88kBq/mL) was closest to the ground truth (843.36843.36843.36~{}843.36kBq/mL), suggesting that algorithms incorporating anatomical information may yield the most accurate kidney activity concentration. For both the simulated and clinical data, the order of algorithms yielding the lowest to highest noise were BSREM-RDP, BSREM-RDPAP, OSEM, and KEM; this is unsurprising as the RDP prior is designed to reduce noise. While BSREM-RDPAP has a lower noise than OSEM in each case, it also had higher uptake in the kidney: BSREM-RDPAP thus has simultaneous improvements in both bias and noise reduction. For the bone lesion, algorithms incorporating anatomical information (BSREM-RDPAP and KEM) were able to contain activity mainly inside the bone; this is demonstrated in the axial slices shown in Figure 3. These algorithms also yielded higher bone lesion activity concentration predictions than OSEM for both the simulated/clinical data, with the simulated data concentrations (299.11299.11299.11~{}299.11kBq/mL for BSREM-RDPAP and 231.30231.30231.30~{}231.30kBq/mL for KEM) closer to the ground truth (819.79819.79819.79~{}819.79kBq/mL) than OSEM (199.07199.07199.07~{}199.07kBq/mL). For the liver lesion in the simulated data, however, use of BSREM-RDPAP actually yielded a smaller activity concentration prediction (244.30244.30244.30~{}244.30kBq/mL) compared to OSEM (251.02251.02251.02~{}251.02kBq/mL). This occured because there was no anatomical difference in the surrounding tissue in the liver metastasis, while there was for the bone metastasis. These results, which are consistent with previous studies [6], suggest that anatomical priors may yield more accurate activity concentrations for bone lesion dosimetry. While this example focused primarily on organ and lesion dosimetry, PyTomography users will be to assess the merits of various reconstruction algorithms on any task specific metric.
@@ Appendix A Validation of PyTomography
    Monte Carlo SPECT Validation With STIR: Radioactivity concentrations in a digital phantom were selected to correspond to a 24-hour post injection prostate cancer patient (1700 MBq total activity) imaged via the prostate-specific membrane antigen (PSMA)-targeting radiopharmaceutical, 177Lu-PSMA-617. SPECT acquisition was simulated in SIMIND with dimensions $128\times 384$, a pixel size of $0.48~{}\text{cm}\times 0.48~{}\text{cm}$, 120 projections, an acquisition time of 15 s/projection, and use of Siemens medium energy collimators. Data were reconstructed with OSEM (2 iterations, 8 subsets) using (i) STIR and (ii) PyTomography. Coronal slices and 1D profiles of reconstructions are shown in Figure 5 and are nearly identical. The image noise, quantified by the standard deviation of predicted counts in the liver, was 0.789 counts (STIR) and 0.793 counts (PyTomography). The mean predicted counts in the kidneys were 1.764 (STIR) and 1.761 (PyTomography).
@@
Figure 7: Sample coronal slices (top) and 1D profiles (bottom) of the high resolution brain phantom reconstructed using CASToR and PyTomography with LM TOF OSEM (2it/21ss). The two images were normalized such that the sum of their voxels was identical. 1D profiles correspond to the horizontal lines shown on the coronal slices. The percent difference in grey and white matter mean uptake between the two software packages were 0.90% and 1.14% respectively.